# Towards Automated Identification of Violation Symptoms of Architecture Erosion

### v4

@@ 3.5. Phase 2 - Validation: Survey and Interview
Data Analysis. For the survey questions, we applied descriptive statistics to analyze the responses to the demographic and Likert scale questions. For the open-ended interview questions, interviews were recorded and transcribed; subsequently, we used the open coding and constant comparison method (Adolph2011CC, ) to analyze and categorize the qualitative responses.
@@ 4.1. RQ1: Identifying Violation Symptoms
To answer RQ1, we trained five ML classifiers and one DL classifier based on three word embedding techniques (i.e., word2vec, fastText, and GloVe). Moreover, we evaluated the performance of these classifiers and compared the difference between their performance.
@@
Table 5. Performance comparison of ensemble classifiers
@@
For ML-based classifiers, as shown in Table 2, we trained the five types of ML classifiers with three word embedding techniques (i.e., word2vec, fastText, and GloVe), leading to 15 classifiers (i.e., the five ML models times three word embeddings). We calculated the average values of the classifiers over the three pre-trained word embeddings. For each classifier based on three word embeddings, we underlined the best result of each metric (horizontal comparison); for each metric of the classifiers (precision, recall, F1, and accuracy), we marked the best metric result of each classifier on three word embeddings in bold (vertical comparison). From the results, we can see that the kNN classifiers have better scores of Precision than other classifiers with an average Precision of 0.816. As for the remaining metrics, the SVM classifiers achieve relatively better performance on nearly all metrics than other classifiers, with an average Recall of 0.801, F1-score of 0.779, and Accuracy of 0.773. Specifically, the SVM classifier based on word2vec performed the best with a Precision of 0.789, a Recall of 0.828, an F1-score of 0.808, and an Accuracy of 0.803.
@@
The ML-based classifiers demonstrated superior performance, achieving, on average, a Precision of 0.753, a Recall of 0.667, an F1-score of 0.700, and an Accuracy of 0.719. In contrast, the DL-based classifiers exhibited lower performance, with an average Precision of 0.574, Recall of 0.664, F1-score of 0.542, and Accuracy of 0.563. These findings suggest that the ML-based classifiers are more effective in accurately classifying the target variable than the DL-based classifiers. Overall, among these classifiers, the SVM classifiers outperform other classifiers, exhibiting the best average results on all metrics, except for the Precision metric.
@@ 4.2. RQ2: Improving Performance with Voting
To answer RQ2, we employed an ensemble learning technique, i.e., majority voting strategy, to build an ensemble classifier. We compared the performance of the individual classifiers with the ensemble classifiers. More specifically, we explored to what extent combining one type of classifier based on three word embedding techniques can achieve better performance, as shown in Table 6 and Table 6.
@@
According to Table 6, when we utilized the voting strategy for each ML classifier based on the three word embeddings (i.e., word2vec, fastText, GloVe), we can see that all ensemble classifiers perform better than the mean values of individual classifiers, with F1-score improvement from 0.48% to 13.10%. Even compared with the best-performing classifiers, most of the ensemble classifiers outperform the individual ML/DL classifiers, except for kNN. Moreover, the ensemble TextCNN of the three TextCNN classifiers based on pre-trained word embeddings (see Table 3) does not perform as well as TextCNN_FT (i.e., the best one) on Precision, F1-score, and Accuracy. This means that TextCNN_FT with the fastText pre-trained word embedding performs best compared with the other three TextCNN classifiers, and the voting strategy does not have an obvious improvement for the trained DL classifiers.
In addition, as shown in Table 6, we utilized the voting strategy for the five ML classifiers based on each of the three word embeddings. For example, ML_word2vec is the ensemble classifier of the five ML classifiers based on the word2vec word embedding (i.e., SVM_w2v, LR_w2v, NB_w2v, DT_w2v, and kNN_w2v). Similarly as before, the “Mean” column refers to the average metric results of ML classifiers (see the last row in Table 2); the “Best” column presents the best-performing classifiers in Table 2; the “Voting” column denotes the performance results of the ensemble classifiers after voting. It is clear that the three ensemble classifiers outperform the average performance results of the individual classifiers, with improvements from 3.11% to 7.98%. Even compared with the best-performing individual classifiers (i.e., the maximum values of the voting samples), the ensemble classifier ML_fastText can still demonstrate performance improvement on all metrics.
Moreover, we can observe the performance difference of ensemble classifiers with three word embeddings. In general, for the selected five ML models, the fastText pre-trained word embedding model can achieve relatively better performance improvements on almost all the metrics compared with the models using word2vec and GloVe. This finding affirms a similar observation in the result of RQ1.2.
RQ2 Summary. In comparison to individual ML classifiers, the ensemble classifiers achieve better performances after utilizing the majority voting strategy. However, no significant improvement was observed in the performance of the DL classifiers on our dataset when implementing the ensemble technique.
@@ 4.3. RQ3: Validation in Practice
As elaborated in Section 3.5, we designed and conducted a survey and semi-structured interviews to solicit the perceptions from the practitioners on the usefulness of our trained models to automatically identify violation symptoms from code review comments.
@@
Overall, we planned to send customized survey invitation emails to 200 code reviewers who were involved in the discussion of the collected review comments concerning architecture violations. However, we could only derive valid email addresses for 169 of them. We finally obtained 24 valid responses to our survey and interviewed 6 participants after sending 169 survey invitation emails and reminder emails; the survey response rate of 14.2%, is a reasonable response rate considering that the general response rate is around 5% in empirical software engineering research (gaese2008CH1, ). Table 7 shows the demographic information of the 24 respondents. Twenty two of the respondents have more than 10 years of experience in software development, while two of them have 6-10 years of experience. Among the respondents, 50.0% (12) are software engineers, 33.3% (8) are developers, and 16.7% (4) are architects. Note that, software engineers usually have a broader work range than developers; for example, the former may work on full-stack development while the latter could deal only with the back end. Regarding the size of their companies, ten of the respondents reported that their companies have more than 1000 employees, five respondents worked in companies with employees between 501 and 1000, six respondents in companies with employees between 101 and 500, and three respondents in companies with employees between 21 and 100. Besides, the respondents worked in a variety of application domains.
Statements 3 to 6 investigate whether the trained classifiers can play an auxiliary role in development and code reviews. Approximately 41.7% respondents strongly agreed (12.5%) or agreed (29.2%) that the identified violation symptoms can help identify violation-related issues faster than manual identification (Statement 4); this indicates that the classifiers may help to speed up code review. Moreover, the majority (62.5%) of the respondents strongly agreed (4.2%) or agreed (58.3%) that automated identification can help locate and prioritize architectural violation issues (Statement 5). Furthermore, 41.7% of the respondents agreed that the identified violation symptoms can inspire them to find out other (or similar) violation symptoms (Statement 6). Notably, one respondent strongly disagreed with all six statements, indicating that he/she might not believe that ML and DL-based classifiers can be useful for identifying architectural violations in practice; we did not treat this response as an outlier. We further analyzed the responses for the six statements and found no significant statistical differences between the identified violation symptoms. Overall, Figure 4 shows that positive feedback (strongly agree + agree) surpasses negative (disagree + strongly disagree) and neural feedback in most of the statements, except for Statement 2.
  • (IQ1) Challenges: Three out of six interviewees raised concerns regarding data leakage problems when employing or integrating such an automated approach supported by third-party libraries or specific Large Language Models (LLMs), especially for inner-source projects. Besides, two out of six interviewees expressed their concerns about false positives; as I1 mentioned: “Any tool that gives you warnings that are false, after a while you start ignoring its warnings”. Regardless of the type of automated techniques employed, rich and high-quality training data on violation symptoms and suitable data processing are necessary to ensure the performance and generality of these approaches. As I2 stated, “Training data is quite important for the accuracy of the results, …, there were always garbage outputs even if the input per se is good”. Therefore, putting more effort into improving data processing could pose another challenge for such automated approaches.
  • (IQ2) Strategies: I4 thought that the best practices and specific rules for identifying typical violation symptoms (e.g., layering violations) should be considered and integrated when employing automated techniques during code review. This could help customize such an approach of identifying violation symptoms and then maximize the identification accuracy for inner-source projects inside companies.
RQ3 Summary. Practitioners’ responses indicate the usefulness of the automated identification of violation symptoms discussed in code review comments in practice. Besides, the identified violation symptoms can enable and inspire practitioners to find architecturally related violations and prioritize violation-related issues.
@@
P2 Germany More than 10 years Developer 21 $\leq$ employees $\leq$ 100 Industrial electronics
@@
P6 Germany 6 - 10 years Developer 101 $\leq$ employees $\leq$ 500 Development tools
P7 USA More than 10 years Developer 501 $\leq$ employees $\leq$ 1000 Iaas
P8 Norway More than 10 years Developer 501 $\leq$ employees $\leq$ 1000 Cross-platform libraries for application development
P9 Norway More than 10 years Developer 101 $\leq$ employees $\leq$ 500 SW development frameworks
@@
P11 Germany More than 10 years Developer 101 $\leq$ employees $\leq$ 500 Middleware
@@
P15 Germany More than 10 years Developer 501 $\leq$ employees $\leq$ 1000 Tools and frameworks for software development
P16 (I2) Germany More than 10 years Developer 101 $\leq$ employees $\leq$ 500 Middleware Platform

### v5

@@ 3.4. Phase 2 - Validation: Survey and Interview
Data Analysis. For the survey questions, we applied descriptive statistics to analyze the responses to the demographic and Likert scale questions. For the open-ended interview questions, interviews were recorded and transcribed; subsequently, we used the Open Coding and Constant Comparison method (Adolph2011CC, ) (a fundamental qualitative analysis technique widely used in empirical software engineering research) to analyze and categorize the qualitative responses.
3.5. Phase 3 - Comparison: LLMs vs. ML/DL Techniques
Unlike traditional ML/DL approaches, LLMs eliminate the need for task-specific training from scratch. Therefore, we leveraged pre-trained and open-source LLMs using techniques like prompt engineering to perform the same classification tasks in Phases 1 and 2. First, we conducted preliminary experiments to identify high-performing prompts and adopted the best-performing prompt for subsequent formal evaluations. We employed both zero-shot and few-shot learning approaches, explicitly evaluating the few-shot setting with 0, 2, 4, 6, and 10 examples, and assessed how performance varied across three representative LLMs: GPT-4o, Qwen-2.5, and DeepSeek-R1.
LLM Selection. We selected three representative and high-performance LLMs based on their technical diversity, capabilities, and availability.
  • GPT-4o: A proprietary multimodal LLM developed by OpenAI (GPT-4o, ). Although its architecture is undisclosed, industry speculation suggests that it involves approximately 200 billion parameters. GPT-4o demonstrates state-of-the-art performance in code generation (HumanEval) and mathematical reasoning (MATH), and supports real-time multimodal interactions with an average response latency of 320 ms. Primary applications include enterprise AI, educational systems, and multimedia content generation.
  • Qwen-2.5: An open-source LLM from Alibaba’s DAMO Academy with 72 billion parameters. Qwen2.5-72B (qwen2.5, ) supports a 128k-token context window and excels in multilingual performance, particularly in Chinese, as demonstrated by the C-Eval benchmark. It supports 12 languages and achieves ChatGPT-4-level comprehension.
  • DeepSeek-R1: An open-source 671-billion-parameter model (deepseekai2025, ) using a Mixture of Experts (MoE) design with only 8 billion active parameters per inference. This architecture offers a 60% reduction in training cost and over 2.3× inference throughput compared to dense models like LLaMA 3-70B. Its efficient MoE architecture and strong reasoning capabilities make it particularly suitable for evaluating performance across diverse SE tasks while maintaining computational efficiency.
Zero-Shot and Few-Shot Learning. LLMs inherently support zero-shot and few-shot learning due to their robust generalization and natural language understanding capabilities.
  • Zero-shot learning involves presenting no labeled examples. Instead, the model receives only a conceptual description (e.g., “violation symptoms”) and must perform classification based on this abstract guidance alone.
  • Few-shot learning supplements the conceptual definition with a small set of labeled examples. The model uses these demonstrations to infer classification strategies, blending prior knowledge with task-specific cues.
To investigate performance variability, we tested four different few-shot settings (2, 4, 6, and 10 examples), ensuring class balance by selecting even numbers of samples.
Prompt Engineering. Prompt engineering is critical in guiding LLMs toward accurate and reliable outputs. Properly designed prompts significantly influence performance, particularly in complex classification tasks. While optimal prompt design is often non-trivial due to the large design space, we adopted an empirical evaluation approach. We generated multiple prompt candidates for both zero-shot and few-shot learning and evaluated their performance in preliminary trials, and then selected the best-performing prompt. The prompt defined the LLM’s assumed role, introduced the concept of violation symptoms, provided demonstrations (in few-shot settings), and specified the required output format.
@@ 4.1. RQ1: Identifying Violation Symptoms
To answer RQ1, we trained five ML classifiers and one DL classifier based on three word embedding techniques (i.e., word2vec, fastText, and GloVe). Moreover, we evaluated the performance of these classifiers and compared the differences between their performance. In general, we use F1-score as the overall evaluation of classifiers’ performance.
@@
Table 5. Performance comparison of ensemble ML/DL-based classifiers
@@
For ML-based classifiers, as shown in Table 2, we trained the five types of ML classifiers with three word embedding techniques (i.e., word2vec (w2v), fastText (FT), and GloVe), leading to 15 classifiers (i.e., the five ML models times three word embeddings). We calculated the average values of the classifiers over the three pre-trained word embeddings. For each classifier based on three word embeddings, we underlined the best result of each metric (horizontal comparison); for each metric of the classifiers (precision, recall, F1-score, and accuracy), we marked the best metric result of each classifier on three word embeddings in bold (vertical comparison). From the results, we can see that the kNN classifiers have better scores of Precision than other classifiers with an average Precision of 0.816. As for the remaining metrics, the SVM classifiers achieve relatively better performance on nearly all metrics than other classifiers, with an average Recall of 0.801, F1-score of 0.779, and Accuracy of 0.773. Specifically, the SVM classifier based on word2vec performed the best with a Precision of 0.789, a Recall of 0.828, an F1-score of 0.808, and an Accuracy of 0.803.
@@
The ML-based classifiers demonstrated superior performance, achieving, on average, a Precision of 0.753, a Recall of 0.667, an F1-score of 0.700, and an Accuracy of 0.719. In contrast, the DL-based classifiers exhibited lower performance, with an average Precision of 0.574, Recall of 0.664, F1-score of 0.542, and Accuracy of 0.563. These findings suggest that the ML-based classifiers are more effective in accurately classifying the target variable than the DL-based classifiers in our case. Overall, among these classifiers, the SVM classifiers outperform other classifiers, exhibiting the best average results on all metrics, except for the Precision metric.
@@ 4.2. RQ2: Improving Performance with Voting
To answer RQ2, we employed an ensemble learning technique, i.e., majority voting strategy, to build an ensemble classifier. We compared the performance of individual classifiers with ensemble classifiers. Specifically, we explored to what extent combining one type of classifier based on three word embedding techniques can achieve better performance, as shown in Table 6 and Table 6.
@@
According to Table 6, when we utilized the voting strategy for each ML classifier based on the three word embeddings, all ensemble classifiers perform better than the mean values of individual classifiers, with F1-score improvement from 0.48% to 13.10%. Table 6 shows that even compared with the best-performing classifiers, most of the ensemble classifiers (4 out of 5) outperform the individual ML/DL classifiers, except for kNN. Moreover, the ensemble TextCNN of the three TextCNN classifiers based on pre-trained word embeddings (see Table 3) does not perform as well as TextCNN_FT (i.e., the best one) on Precision, F1-score, and Accuracy. This means that TextCNN_FT with the fastText pre-trained word embedding performs best compared with the other three TextCNN classifiers, and the voting strategy does not have an obvious improvement for the trained DL classifiers.
Additionally, as shown in Table 6, we utilized the voting strategy for the five ML classifiers based on each of the three word embeddings. For example, ML_word2vec is the ensemble classifier of the five ML classifiers based on the word2vec word embedding (i.e., SVM_w2v, LR_w2v, NB_w2v, DT_w2v, and kNN_w2v). Similarly as before, the “Mean” column refers to the average metric results of ML classifiers (see the last row in Table 2); the “Best” column presents the best-performing classifiers in Table 2; the “Voting” column denotes the performance results of the ensemble classifiers after voting. It is clear that the three ensemble classifiers outperform the average performance results of the individual classifiers, with improvements from 3.11% to 7.98%. Even compared with the best-performing individual classifiers (i.e., the maximum values of the voting samples), the ensemble classifier ML_fastText can still demonstrate performance improvement on all metrics.
In general, regarding the performance difference of ensemble classifiers with three word embeddings, among the selected five ML models, the fastText pre-trained word embedding model achieves relatively better performance improvements on almost all metrics compared with the models using word2vec and GloVe. This finding affirms a similar observation in the result of RQ1.2.
RQ2 Summary. In comparison to individual ML classifiers, the ensemble classifiers achieve better performances after utilizing the majority voting strategy. However, no significant improvement was observed in the performance of the DL classifiers on our dataset when implementing the ensemble strategy.
@@ 4.3. RQ3: Validation in Practice
As elaborated in Section 3.4, we designed and conducted a survey and semi-structured interviews to solicit the perceptions from the practitioners on the usefulness of our trained models to automatically identify violation symptoms from code review comments. Note that, our survey and interview did not ask participants to “discover” additional violation symptoms that were not originally discussed in their code reviews.
@@
Overall, we planned to send customized survey invitation emails to 200 code reviewers who were involved in the discussion of the collected review comments concerning architecture violations. However, we could only derive valid email addresses for 169 of them. We finally obtained 24 valid responses to our survey and interviewed 6 participants after sending 169 survey invitation emails and reminder emails; the survey response rate of 14.2%, is a reasonable response rate considering that the general response rate is around 5% in empirical software engineering research (gaese2008CH1, ). Table 7 shows the demographic information of the 24 respondents. Twenty two of the respondents have more than 10 years of experience in software development, while two of them have 6-10 years of experience. Among the respondents, 83.3% (20) are software engineers and 16.7% (4) are architects. Regarding the size of their companies, ten of the respondents reported that their companies have more than 1000 employees, five respondents worked in companies with employees between 501 and 1000, six respondents in companies with employees between 101 and 500, and three respondents in companies with employees between 21 and 100. Besides, the respondents worked in a variety of application domains.
@@
P2 Germany More than 10 years Software Engineer 21 $\leq$ employees $\leq$ 100 Industrial electronics
@@
P6 Germany 6 - 10 years Software Engineer 101 $\leq$ employees $\leq$ 500 Development tools
P7 USA More than 10 years Software Engineer 501 $\leq$ employees $\leq$ 1000 Iaas
P8 Norway More than 10 years Software Engineer 501 $\leq$ employees $\leq$ 1000 Cross-platform libraries for application development
P9 Norway More than 10 years Software Engineer 101 $\leq$ employees $\leq$ 500 SW development frameworks
@@
P11 Germany More than 10 years Software Engineer 101 $\leq$ employees $\leq$ 500 Middleware
@@
P15 Germany More than 10 years Software Engineer 501 $\leq$ employees $\leq$ 1000 Tools and frameworks for software development
P16 (I2) Germany More than 10 years Software Engineer 101 $\leq$ employees $\leq$ 500 Middleware Platform
@@
Statements 3 to 6 investigate whether the trained classifiers can play an auxiliary role in development and code reviews. Approximately 41.7% respondents strongly agreed (12.5%) or agreed (29.2%) that the identified violation symptoms can help identify violation-related issues faster than manual identification (Statement 4), while there are still nearly half of the participants convey their uncertainty; the proportion of positive responses indicates that the classifiers may not fully replace manual efforts, but to some extent they can speed up the identification of violation-related issues during code review. Moreover, the majority (62.5%) of the respondents strongly agreed (4.2%) or agreed (58.3%) that automated identification can help locate and prioritize architectural violation issues (Statement 5). Furthermore, 41.7% of the respondents agreed that the identified violation symptoms can inspire them to find out other (or similar) violation symptoms (Statement 6). Notably, one respondent strongly disagreed with all six statements, indicating that he/she might not believe that ML/DL-based classifiers can be useful for identifying architectural violations in practice; we did not treat this response as an outlier. Moreover, only a small number of participants (ranging from 2 for Statement 1 to 5 for Statement 6) chose the “I Don’t Know” option. We further analyzed the responses for the six statements and found no significant statistical differences between the identified violation symptoms. Overall, Figure 4 shows that positive feedback (strongly agree + agree) surpasses negative (disagree + strongly disagree) and neutral feedback in most of the statements, except for Statement 2.
  • (IQ1) Challenges: Three out of six interviewees raised concerns regarding data leakage problems when employing or integrating such an automated approach supported by third-party libraries or specific LLMs, especially for inner-source projects. Besides, two out of six interviewees expressed their concerns about “Crying Wolf” effect of false positives; as I1 mentioned: “Any tool that gives you warnings that are false, after a while you start ignoring its warnings”. If tools give false warnings (or warnings that users consider trivial), it eventually leads users to ignore all warnings, including legitimate ones. This calls for the interpretability of such automated approaches that participants might not only want to know about the identified violation symptoms, but also understand why they are flagged as “violations”; as I2 mentioned: “As a user, I am not satisfied to be told where the issues are located, and I also want to know why they are flagged as violations with explanations”. Therefore, putting more effort into improving the interpretability could pose another challenge for such automated approaches.
  • (IQ2) Strategies: I4 thought that the best practices and specific rules (e.g., guidelines on architectural layering, module dependencies, and coupling thresholds) for identifying typical violation symptoms (e.g., layering violations) should be considered and integrated when employing automated techniques during code review. This could help customize such an approach of identifying violation symptoms and then maximize the identification accuracy for inner-source projects inside companies.
RQ3 Summary. Practitioners’ responses indicate that the automated identification of violation symptoms is perceived as useful in practice. In particular, it helps practitioners identify potential violation symptoms earlier, better recognize architecturally related violations and prioritize violation-related issues during review and maintenance activities.
4.4. RQ4: Comparison with LLM-based Classifiers
To answer RQ4, we conducted additional comparative experiments by constructing LLM-based classifiers. As shown in Table 8, we evaluated three LLMs across five shot settings (one zero-shot and four few-shot configurations). Qwen-2.5 demonstrates the highest Precision, achieving a peak of 0.873 in the 4-shot setting and the highest average Precision of 0.839. In contrast, GPT-4o excels in Recall, with a maximum value of 0.951 at 10-shot and an average Recall of 0.879. DeepSeek-R1 also performs well in Recall, particularly at 4-shot (0.943), resulting in the highest average Recall of 0.898. For F1-scores, GPT-4o achieves the highest peak 0.851 at 2-shot and the highest average 0.811, followed by Qwen-2.5 (0.782 at 6-shot) and DeepSeek-R1 (0.767 at 4-shot). GPT-4o also leads in Accuracy, reaching a peak of 0.848 at 2-shot and averaging 0.793. Qwen-2.5 achieves its highest Accuracy (0.799) at 6-shot, while DeepSeek-R1 has the lowest average Accuracy (0.696) despite a 2-shot peak of 0.713.
Overall, GPT-4o provides the most balanced and robust performance across all metrics and shot settings. Qwen-2.5 leads in Precision and maintained competitive Accuracy, while DeepSeek-R1 stands out in recall but shows lower Precision and overall Accuracy.
Table 8. Performance comparison of different LLMs with varying shot settings
Model Precision Recall
0-shot 2-shot 4-shot 6-shot 10-shot Average 0-shot 2-shot 4-shot 6-shot 10-shot Average
Qwen-2.5 0.825 0.831 0.873 0.854 0.811 0.839 0.656 0.566 0.566 0.721 0.705 0.643
GPT-4o 0.701 0.835 0.851 0.699 0.712 0.759 0.844 0.869 0.795 0.934 0.951 0.879
Deepseek-R1 0.653 0.657 0.646 0.611 0.635 0.640 0.803 0.893 0.943 0.926 0.926 0.898
Average 0.726 0.774 0.790 0.721 0.719 0.746 0.768 0.776 0.768 0.860 0.861 0.807
Model F1-score Accuracy
0-shot 2-shot 4-shot 6-shot 10-shot Average 0-shot 2-shot 4-shot 6-shot 10-shot Average
Qwen-2.5 0.731 0.673 0.687 0.782 0.754 0.725 0.758 0.725 0.742 0.799 0.771 0.759
GPT-4o 0.766 0.851 0.822 0.800 0.814 0.811 0.742 0.848 0.828 0.766 0.783 0.793
Deepseek-R1 0.721 0.757 0.767 0.736 0.753 0.747 0.689 0.713 0.713 0.668 0.697 0.696
Average 0.739 0.760 0.759 0.773 0.774 0.761 0.730 0.762 0.761 0.744 0.750 0.749
Compared to ML/DL-based classifiers (see Tables 2 and  3), the LLM-based classifiers achieve better performance on the same classification task, with average Precision of 0.746, Recall of 0.807, F1-score of 0.761, and Accuracy of 0.749. This finding indicates that the LLM-based classifiers are more effective in identifying the violation symptoms than traditional ML/DL-based classifiers. Among the LLM-based classifiers, the GPT-4o model with a 2-shot setting outperforms others in average performance across all metrics except Recall.
We also examined whether ensemble strategy across different shot settings could improve performance (see Tables 10 and 10). Table 10 shows the results of ensemble classifiers based on a voting strategy. The “Mean” column refers to the average performance of each metric across all LLMs and shot settings from Table 8, and the “Best” column captures the best performance per shot setting. The “Voting” column reports the ensemble performance obtained via majority voting among models with the same shot setting. Likewise, the improvements over the “Mean” and “Best” are shown in the “Impro_M” and “Impro_B” columns, respectively. We observed that the ensemble classifiers outperform the average results of different shot settings across all metrics in Table 10. However, they generally underperform when compared to the best individual classifiers in Precision and Recall. In contrast, the ensemble classifiers slightly improve over the best individual classifiers in F1-score and Accuracy.
Similarly, Table 10 examines ensemble classifiers within a single LLM across different shot settings. The “Mean” column presents the average metric values across all five shot settings from Table 8; the “Best” column shows the best performance of single LLMs; the “Voting” column displays the results from the ensemble classifiers across shot settings using a voting strategy. Table 10 shows that, for GPT-4o and Qwen-2.5, the ensemble classifiers outperform their respective average performances but do not exceed their best individual results. Notably, for Deepseek-R1, the ensemble classifier surpasses both the average and best individual classifiers across all metrics except Precision.
Table 9. Performance comparison of ensemble LLM-based classifiers with different shot settings
Classifier Precision Recall F1-score Accuracy
0-shot 0.726 0.825 0.760 4.65% -7.85% 0.768 0.844 0.779 1.42% -7.77% 0.739 0.766 0.769 4.09% 0.45% 0.730 0.758 0.766 5.06% 1.08%
2-shot 0.774 0.835 0.813 5.01% -2.59% 0.776 0.893 0.820 5.63% -8.26% 0.761 0.851 0.816 7.34% -4.12% 0.762 0.848 0.816 6.99% -3.86%
4-shot 0.790 0.873 0.845 6.92% -3.27% 0.768 0.943 0.803 4.63% -14.78% 0.758 0.822 0.824 8.58% 0.18% 0.761 0.828 0.828 8.80% 0.00%
6-shot 0.722 0.854 0.717 -0.61% -16.07% 0.861 0.934 0.893 3.81% -4.39% 0.773 0.800 0.796 2.95% -0.55% 0.745 0.799 0.770 3.49% -3.59%
10-shot 0.719 0.811 0.723 0.46% -10.94% 0.861 0.951 0.918 6.67% -3.45% 0.774 0.814 0.809 4.49% -0.66% 0.750 0.783 0.783 4.37% 0.00%
Table 10. Performance comparison among ensemble LLM-based classifiers
Classifier Precision Recall F1-score Accuracy
Qwen-2.5 0.839 0.873 0.849 1.24% -2.74% 0.643 0.721 0.648 0.77% -10.23% 0.725 0.782 0.735 1.31% -6.05% 0.759 0.799 0.766 0.97% -4.10%
GPT-4o 0.759 0.851 0.803 5.72% -5.64% 0.879 0.951 0.902 2.61% -5.17% 0.811 0.851 0.849 4.78% -0.23% 0.793 0.848 0.840 5.89% -0.97%
Deepseek-R1 0.640 0.657 0.650 1.47% -1.05% 0.898 0.943 0.943 4.93% 0.00% 0.747 0.767 0.769 3.01% 0.33% 0.696 0.713 0.717 3.06% 0.57%
RQ4 Summary. GPT-4o shows the most balanced performance, leading in F1-score and Accuracy, while Qwen-2.5 excels in Precision and DeepSeek-R1 in Recall. LLM-based classifiers outperform traditional ML/DL-based classifiers across all metrics. Ensemble classifiers offer slight improvements in F1-score and Accuracy but do not surpass the best individual LLM-based classifiers in Precision and Recall.